# Adversarial Disentanglement by Backpropagation with Physics-Informed Variational Autoencoder

### v2

@@ 1 Introduction
Our objective is to accurately infer a posterior distribution over physically meaningful latent variables, to reconstruct the structural response, quantify the associated uncertainty, and predict the damage and degradation condition of the system in previously unseen conditions. This is achieved using a limited number of noisy measurements of the structural response, domain and class variables. Due to the influence of the domain, class, and other unknown confounding factors, this will generally be an ill-posed inverse problem that requires learning a disentangled representation [Bengio2014] of the different generative factors. This task is further complicated by the limitations of physics-based models, which often represent structures under idealized nominal conditions and disregard the influence of environmental and operational variability, damage, and degradation. Most computational models of physical systems will contain simplifications and approximations due to lack of knowledge about certain aspects of the underlying physical process and to ensure computational tractability. Reducing this epistemic uncertainty is often infeasible due to cost or time constraints. As a result, only a partial description of the physical system is available in practical applications.
@@
Generative probabilistic models such as Variational Autoencoders (VAE) [Kingma2022], Normalizing Flows [Rezende2016], and Generative Adversarial Networks (GANs) [Goodfellow2014], are a class of models that employ deep learning architectures to approximate the distribution of a given set of data and generate samples from the learned distribution. Generative probabilistic models have recently seen broader use in Structural Health Monitoring (SHM), and for constructing digital twins of structures [Bacsa2025, Coraca2023, Tsialiamanis2021, Mao2021]. We propose a VAE architecture for approximating the joint distribution between the structural response and a set of physics-grounded latent variables, while accounting and correcting for the confounding influence of the domain and class of the structure, by leveraging observed domain and class variables. To achieve this, the VAE components are split into physics-based and data-driven branches, trained simultaneously in an end-to-end fashion. The data-driven branches are tasked with extracting features of the response that are informative about the domain and class variables, encoding them into the corresponding latent space, and using the latent code to augment the physics-based model predictions. Formulating the VAE as a combination of physics-based and data-driven components is not a straightforward task. The flexibility and learning capacity of feed-forward Neural Networks (NNs) that enables them to accurately model physical processes from data can be problematic when combining them with physics-based models, as the flexible NN components tend to override the known physics [Takeishi2021], resulting in inaccurate inference and overconfident or unrealistic predictions. To address this issue we propose an adversarial training objective that encourages an interpretable and parsimonious representation of the physical system by constraining the data-driven components of the VAE. Once trained, the model can be used to simultaneously perform inference over physically meaningful latent variables for new measurements as they become available, and generate samples from the predictive distribution of the response. Given a set of response measurements, the trained model can also be used to predict the corresponding domain and class variables. The proposed approach aims to:
@@ 2.2 Epistemic Uncertainty in Bayesian Model Updating
Uncertainty in the modeling of physical systems can generally be classified as either aleatoric or epistemic. Aleatoric uncertainty is the component of the uncertainty due to the inherent randomness of a physical process that can not be reduced, while epistemic uncertainty stems from lack of knowledge regarding the physical process [Kiureghian2009]. The reader is referred to [Kamariotis2024] for an extended overview on classification and treatment of uncertainties in SHM applications. Epistemic uncertainty is always present to some degree in practical applications, either due to lack of knowledge or due to simplifications and approximations used to make the evaluation of the physics-based model computationally tractable. Suppose that an analytical or numerical physics-based model of a structure, defined as a function $f(\bm{z}_{\mathrm{x}})$ is available. The response measurements $\bm{x}$ can then be expressed as $\bm{x}=f(\bm{z}_{\mathrm{x}})+\epsilon$, where $\epsilon$ is a realization of a random variable quantifying the discrepancy between $f(\bm{z}_{\mathrm{x}})$ and $\bm{x}$ due to the combined influence of aleatoric and epistemic uncertainties. In the Bayesian model updating framework, the measured response of a physical system is used to update the prior knowledge, expressed in terms of a prior distribution $p(\bm{z}_{\mathrm{x}})$ over physically meaningful latent variables $\bm{z}_{\mathrm{x}}$. This is achieved by approximating the data generating process as a combination of a deterministic physics-based model and a probabilistic model [Kennedy2001], where the latter accounts for the combined influence of epistemic and aleatoric uncertainty. In practical applications, the available domain knowledge is often not sufficient to guarantee that the coupled probabilistic-physical model is an accurate description of the data generating process, i.e. the real-world process that generated the observations. In this work, it is assumed that the epistemic uncertainty stems from the confounding influence of the domain and class of a structure, and our inability (e.g. due to cost or time constraints) to account for these influences in the form and parameters of the physics-based model.
@@ 2.3 The Variational Autoencoder
To remedy the lack of domain knowledge, data-driven models based on machine learning techniques have emerged as an alternative to physics-based models, where the unknown physical process is learned from measurements using flexible parametrized approximations. VAE [Kingma2022, Kingma2019] are a popular data-driven approach for learning a joint distribution of data and the latent variables that are assumed to have generated the data using amortized Variational Inference (VI) [Blei2017]. In VAE, the per-datapoint posterior distribution is approximated using a parametrized family of distributions, where the optimal parameters are obtained by minimizing the Kullback-Leibler divergence (KLD) between the true and approximate posteriors. The VAE is composed of an encoder network $q_{\mathrm{\bm{\phi}}}(\bm{z}|\bm{x})$ and a decoder network $p_{\theta}(\bm{x}|\bm{z})$, parametrized by $\phi$ and $\theta$ respectively, where $\bm{z}$ denotes latent variables that can not be observed directly and must be inferred from measurements. The encoder is typically implemented as a feed-forward NN that maps the inputs $\bm{x}$ to a conditional density over latent variables $\bm{z}$. The decoder network $p_{\mathrm{\bm{\theta}}}(\bm{x}|\bm{z})$ works in the opposite direction by approximating the density of $\bm{x}$ conditioned on $\bm{z}$. The training process for VAE consists of simultaneously optimizing the parameters of the decoder that reconstructs the observations samples of the latent variables, and the encoder that maps inputs to a posterior distribution over these latent variables. Optimization is performed by maximizing a lower bound on the marginal likelihood of the data known as the Evidence Lower BOund (ELBO), denoted as $\mathcal{L}_{\mathrm{VAE}}$ in Equation 2.1. Sampling $z\sim q_{\phi}(\bm{z}|\bm{x})$ and evaluating the decoder yields samples from the learned distribution of the data, which in the context of civil and mechanical structural systems can be used for downstream tasks such as remaining useful life assessment.
@@
While data-driven approaches might excel in accurately predicting the response of a physical system for a given set of input parameters when sufficient training data is available, the resulting models are typically black boxes that lack interpretability and yield no useful insights about the underlying physical process that generated the measurements. In cases where both the domain knowledge and the available data are limited, neither approach is feasible, necessitating a compromise between the two extremes. Physics-enhanced machine learning (PEML) encompasses a wide range of approaches that combine machine learning with domain knowledge [Cicirello2024, HaywoodAlexander2024, Cross2022, Rueden2023]. In this paradigm, the available domain knowledge can be supplemented with data, resulting in more accurate and interpretable models than would be possible with either domain knowledge or data alone. PEML approaches have the potential to reduce the required amount of data, improve accuracy and generalization performance and ensure that model predictions are consistent with the known physics. Importantly, incorporating the known physics can yield interpretable representations of physically meaningful quantities, and models that are robust and explainable.
@@ 2.4 Challenges in combining physics-based and data-driven components in VAE
A straightforward approach to account for epistemic uncertainty in a data-driven manner would be to approximate the measured response $\bm{x}$ as the sum of the physics-based model [^1] It is assumed that the gradients of the function with respect to the inputs can be evaluated efficiently to obtain a computationally tractable optimization problem. $f(\bm{z}_{\mathrm{x}})$ and a trainable NN-based function $g_{\bm{\theta}}(\cdot)$, where the latter corrects the discrepancies between the physics-based model predictions and measurements. This type of hybrid model is referred to as a residual model. However, parametrizing the data driven component of the residual model as $g_{\bm{\theta}}(\bm{z}_{\mathrm{x}})$ is not feasible when it is required that $\bm{z}_{\mathrm{x}}$ is interpretable: The resulting hybrid generative model has a posterior distribution $p_{\mathrm{\theta}}(\bm{z}_{\mathrm{x}}|\bm{x})$, where the latent variables $\bm{z}_{\mathrm{x}}$ are the input to a coupled physics-based and data-driven model, and thus no longer physically meaningful. Instead, the latent space can be partitioned as $(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})\in\bm{z}$ and the data driven component parametrized as $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$, where $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ are physically meaningless latent variables intended to capture variability in the measured response due to the influence of the domain and the class. Assuming that the remaining aleatory uncertainties (e.g. caused by measurement noise) are independent of the signal being measured and can be sufficiently modeled as independent and identically distributed (i.i.d.) samples of Gaussian white noise with standard deviation $\sigma_{\mathrm{x}}$, the response measurements can be expressed as:
@@
$p_{\theta}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}}):=\mathcal{N}\left(\hat{\bm{x}}_{\mathrm{p}}+\hat{\bm{x}}_{\mathrm{d}},\ \sigma^{2}_{\mathrm{x}}\bm{I}\right),$ (2.3)
@@
The observables $\bm{c}$ and $\bm{y}$ can be used to ensure that the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ encode information about the domain and class of the structure by simultaneously training auxiliary tasks $r_{\mathrm{c}}(\bm{c}|\bm{z}_{\mathrm{c}})$ and $r_{\mathrm{y}}(\bm{y}|\bm{z}_{\mathrm{y}})$. The resulting hybrid generative model $p_{\bm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ can be coupled with variational posteriors $q_{\mathrm{\phi}_{\mathrm{x}}}(\bm{z}_{\mathrm{x}}|\bm{x})q_{\mathrm{\phi}_{\mathrm{c}}}(\bm{z}_{\mathrm{c}}|\bm{x})q_{\mathrm{\phi}_{\mathrm{y}}}(\bm{z}_{\mathrm{y}}|\bm{x})$ to yield an architecture similar to VAE. This is the architecture derived in Section 3, without the additional constraints and regularization. This parametrization circumvents the previous issue, but does not ensure that the known physics are utilized by the model, or that $\bm{z}_{\mathrm{x}}$ is physically meaningful. Without further regularization, the model can learn combinations of arbitrary predictions from the physics-based and data-driven components $f(\bm{z}_{\mathrm{x}})$ and $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ that sum to an accurate prediction. To see why, it is sufficient to consider the form of the objective given in Equation 2.1. Both the encoder and decoder are aligned in the task of maximizing the reconstruction term $\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}|\bm{z})\right]$, and the data-driven components of the model will account for discrepancies between the physics-based model prediction and measurements up to some level of noise. This results in an entangled representation, where the data-driven components override the known physics and the physics-grounded latent variables loose their physical meaning. This issue is illustrated in Figure 2 using the beam case study shown in Figure 1(a). Further details of the case study are provided in Section 5.1. In this example, a VAE trained on a dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{c}_{i},\bm{y}_{i})\}_{i=1}^{N}$ is evaluated on a new set of input measurements $\bm{x}$, generated from the ground truth data generating process by linearly varying the position of the load $x_{\mathrm{F}}$. It can be seen that the effect of this variation on the measured response is largely captured by the data-driven component of the decoder, which overrides the known physics, despite the fact that the physics-based model includes the load position as an input parameter. The extent to which the data-driven components override the known physics can be inconsistent and hard to predict, and will depend on the neural network architectures and the physics of the problem at hand.
@@
The results presented in Figure 2 are obtained under the assumption of a factorized variational posterior. Ideally, a single encoder with shared parameters $q_{\mathrm{\phi}}(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}}|\bm{x})$ would be used for the three subsets of the latent space $\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. However, using a shared encoder leads to further degradation of the performance as noted by [Ilse2020]. This example demonstrates how the interaction between the physics-based and data-driven components, which determines the resulting learned representation, depends on the capacity and flexibility of the individual machine learning components. Standard VAE offer no mechanism to control this interaction, and are therefore unable to guarantee that the physics will be utilized correctly in the presence of flexible NN-based decoder components.

### v3

@@ 1 Introduction
Our objective is to accurately infer a posterior distribution over physically meaningful latent variables, to reconstruct the structural response, quantify the associated uncertainty, and predict the damage and degradation condition of the system in previously unseen conditions. This is achieved using a limited number of noisy measurements of the structural response, domain and class variables. Due to the influence of the domain, class, and other unknown confounding factors, this will generally be an ill-posed inverse problem that requires learning a disentangled representation [6] of the different generative factors. This task is further complicated by the limitations of physics-based models, which often represent structures under idealized nominal conditions and disregard the influence of environmental and operational variability, damage, and degradation. Most computational models of physical systems will contain simplifications and approximations due to lack of knowledge about certain aspects of the underlying physical process and to ensure computational tractability. Reducing this epistemic uncertainty is often infeasible due to cost or time constraints. As a result, only a partial description of the physical system is available in practical applications.
@@
Generative probabilistic models such as Variational Autoencoders (VAE) [41], Normalizing Flows [56], and Generative Adversarial Networks (GANs) [29], are a class of models that employ deep learning architectures to approximate the distribution of a given set of data and generate samples from the learned distribution. Generative probabilistic models have recently seen broader use in Structural Health Monitoring (SHM), and for constructing digital twins of structures [5, 14, 68, 50]. We propose a VAE architecture for approximating the joint distribution between the structural response and a set of physics-grounded latent variables, while accounting and correcting for the confounding influence of the domain and class of the structure, by leveraging observed domain and class variables. To achieve this, the VAE components are split into physics-based and data-driven branches, trained simultaneously in an end-to-end fashion. The data-driven branches are tasked with extracting features of the response that are informative about the domain and class variables, encoding them into the corresponding latent space, and using the latent code to augment the physics-based model predictions. Formulating the VAE as a combination of physics-based and data-driven components is not a straightforward task. The flexibility and learning capacity of feed-forward Neural Networks (NNs) that enables them to accurately model physical processes from data can be problematic when combining them with physics-based models, as the flexible NN components tend to override the known physics [63], resulting in inaccurate inference and overconfident or unrealistic predictions. To address this issue we propose an adversarial training objective that encourages an interpretable and parsimonious representation of the physical system by constraining the data-driven components of the VAE. Once trained, the model can be used to simultaneously perform inference over physically meaningful latent variables for new measurements as they become available, and generate samples from the predictive distribution of the response. Given a set of response measurements, the trained model can also be used to predict the corresponding domain and class variables. The proposed approach aims to:
@@ 2.2 Epistemic Uncertainty in Bayesian Model Updating
Uncertainty in the modeling of physical systems can generally be classified as either aleatoric or epistemic. Aleatoric uncertainty is the component of the uncertainty due to the inherent randomness of a physical process that can not be reduced, while epistemic uncertainty stems from lack of knowledge regarding the physical process [42]. Epistemic uncertainty is always present to some degree in practical applications, either due to lack of knowledge or due to simplifications and approximations used to make the evaluation of the physics-based model computationally tractable. The reader is referred to [36] for an extended overview on classification and treatment of uncertainties in SHM applications.
Suppose that an analytical or numerical physics-based model of a structure, defined as a function $f(\bm{z}_{\mathrm{x}})$ is available. The response measurements $\bm{x}$ can then be expressed as $\bm{x}=f(\bm{z}_{\mathrm{x}})+\epsilon$, where $\epsilon$ is a realization of a random variable quantifying the discrepancy between $f(\bm{z}_{\mathrm{x}})$ and $\bm{x}$ due to the combined influence of aleatoric and epistemic uncertainties. In the Bayesian model updating framework, the measured response of a physical system is used to update the prior knowledge, expressed in terms of a prior distribution $p(\bm{z}_{\mathrm{x}})$ over physically meaningful latent variables $\bm{z}_{\mathrm{x}}$. This is achieved by approximating the data generating process (i.e. the real-world process that generated the observations) as a combination of a deterministic physics-based model and a probabilistic model [37], where the latter accounts for the combined influence of epistemic and aleatoric uncertainty. In this work, it is assumed that the epistemic uncertainty stems from the confounding influence of the domain and class of a structure, and our inability (e.g. due to cost or time constraints) to account for these influences in the form and parameters of the physics-based model.
@@ 2.3 The Variational Autoencoder
In practical applications, the available domain knowledge is often not sufficient to guarantee that the coupled probabilistic-physical model is an accurate description of the data generating process, limiting the applicability of physics-based modeling. To remedy the lack of domain knowledge, data-driven models based on machine learning techniques have emerged as an alternative to physics-based models, where the unknown physical process is learned from measurements using flexible parametrized approximations. VAE [41, 40] are a popular data-driven approach for learning a joint distribution of data and the latent variables that are assumed to have generated the data using amortized Variational Inference (VI) [7]. In VAE, the per-datapoint posterior distribution is approximated using a parametrized family of distributions, where the optimal parameters are obtained by minimizing the Kullback-Leibler divergence (KLD) between the true and approximate posteriors. The VAE is composed of an encoder network $q_{\mathrm{\bm{\phi}}}(\bm{z}|\bm{x})$ and a decoder network $p_{\theta}(\bm{x}|\bm{z})$, parametrized by $\phi$ and $\theta$ respectively, where $\bm{z}$ denotes latent variables that can not be observed directly and must be inferred from measurements. The encoder is typically implemented as a feed-forward NN that maps the inputs $\bm{x}$ to a conditional density over latent variables $\bm{z}$. The decoder network $p_{\mathrm{\bm{\theta}}}(\bm{x}|\bm{z})$ works in the opposite direction by approximating the density of $\bm{x}$ conditioned on $\bm{z}$. The training process for VAE consists of simultaneously optimizing the parameters of the decoder that reconstructs the observations given samples of the latent variables, and the encoder that maps inputs to a posterior distribution over these latent variables. Optimization is performed by maximizing a lower bound on the marginal likelihood of the data known as the Evidence Lower BOund (ELBO), denoted as $\mathcal{L}_{\mathrm{VAE}}$ in Equation 2.1. Sampling $z\sim q_{\phi}(\bm{z}|\bm{x})$ and evaluating the decoder yields samples from the learned distribution of the data, which in the context of civil and mechanical structural systems can be used for downstream tasks such as remaining useful life assessment.
@@
While data-driven approaches might excel in accurately predicting the response of a physical system for a given set of input parameters when sufficient training data is available, the resulting models are typically black boxes that lack interpretability and yield no useful insights about the underlying physical process that generated the measurements. In cases where both the domain knowledge and the available data are limited, purely physics-based or data-driven approaches become infeasible, necessitating a compromise between the two extremes. Physics-enhanced machine learning (PEML) encompasses a wide range of approaches that combine machine learning with domain knowledge [13, 31, 16, 58]. In this paradigm, the available domain knowledge can be supplemented with data, resulting in more accurate and interpretable models than would be possible with either domain knowledge or data alone. PEML approaches have the potential to reduce the required amount of data, improve accuracy and generalization performance and ensure that model predictions are consistent with the known physics. Importantly, incorporating the known physics can yield interpretable representations of physically meaningful quantities, and models that are robust and explainable.
@@ 2.4 Challenges in combining physics-based and data-driven components in VAE
A straightforward approach to account for epistemic uncertainty in a data-driven manner would be to approximate the measured response $\bm{x}$ as the sum of the physics-based model [^1] It is assumed that the gradients of the function with respect to the inputs can be evaluated efficiently to obtain a computationally tractable optimization problem. $f(\bm{z}_{\mathrm{x}})$ and a trainable NN-based function $g_{\bm{\theta}}(\cdot)$, where the latter corrects the discrepancies between the physics-based model predictions and measurements. This type of hybrid model is referred to as a residual model. Parametrizing the data driven component of the residual model as $g_{\bm{\theta}}(\bm{z}_{\mathrm{x}})$ is not feasible when it is required that $\bm{z}_{\mathrm{x}}$ is interpretable: The resulting hybrid generative model has a posterior distribution $p_{\mathrm{\theta}}(\bm{z}_{\mathrm{x}}|\bm{x})$, where the latent variables $\bm{z}_{\mathrm{x}}$ are the input to a coupled physics-based and data-driven model, and thus no longer physically meaningful. Instead, the latent space can be partitioned as $(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})\in\bm{z}$ and the data driven component parametrized as $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$, where $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ are physically meaningless latent variables intended to capture variability in the measured response due to the influence of the domain and the class. Assuming that the remaining aleatory uncertainties (e.g. caused by measurement noise) are independent of the signal being measured and can be sufficiently modeled as independent and identically distributed (i.i.d.) samples of Gaussian white noise with standard deviation $\sigma_{\mathrm{x}}$, the response measurements can be expressed as:
@@
$p_{\theta}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})\mathrel{\mathop{\ordinarycolon}}=\mathcal{N}\left(\hat{\bm{x}}_{\mathrm{p}}+\hat{\bm{x}}_{\mathrm{d}},\ \sigma^{2}_{\mathrm{x}}\bm{I}\right)$ (2.3)
The observables $\bm{c}$ and $\bm{y}$ can be used to ensure that the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ encode information about the domain and class of the structure by simultaneously training auxiliary tasks $r_{\mathrm{c}}(\bm{c}|\bm{z}_{\mathrm{c}})$ and $r_{\mathrm{y}}(\bm{y}|\bm{z}_{\mathrm{y}})$. The resulting hybrid generative model $p_{\bm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ can be coupled with variational posteriors $q_{\mathrm{\phi}_{\mathrm{x}}}(\bm{z}_{\mathrm{x}}|\bm{x})q_{\mathrm{\phi}_{\mathrm{c}}}(\bm{z}_{\mathrm{c}}|\bm{x})q_{\mathrm{\phi}_{\mathrm{y}}}(\bm{z}_{\mathrm{y}}|\bm{x})$ to yield an architecture similar to VAE. This is the architecture derived in Section 3, without the additional constraints. It should be noted that the assumption of an additive structure for the discrepancy term $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ and the uncertainty term $\bm{\epsilon}_{\mathrm{x}}$ presented in Equation 2.2 is suitable for many physical systems and is commonly employed in hybrid models [16]. We use it without loss of generality with the aim of promoting clarity and interpretability. Depending on domain knowledge regarding the problem at hand, a multiplicative or other form can also be specified. The implications of the additivity assumption are further discussed in Sections 3.2.3 and 6.2.
@@
Neither the additive structure of the hybrid physics-based and data-driven model, nor the specified parametrization of the residual term $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ ensure that the known physics will be utilized by the model, or that $\bm{z}_{\mathrm{x}}$ will be physically meaningful. Without further constraints, the model can learn combinations of arbitrary predictions from the physics-based and data-driven components $f(\bm{z}_{\mathrm{x}})$ and $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ that sum to an accurate prediction. To see why, it is sufficient to consider the form of the objective given in Equation 2.1. Both the encoder and decoder are aligned in the task of maximizing the reconstruction term $\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}|\bm{z})\right]$, and the data-driven components of the model will account for discrepancies between the physics-based model prediction and measurements up to some level of noise. This results in an entangled representation, where the data-driven components override the known physics and the physics-grounded latent variables loose their physical meaning.
This issue is illustrated in Figure 2 using the beam case study shown in Figure 1(a). Further details of the case study are provided in Section 5.1. In this example, a VAE trained on a dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{c}_{i},\bm{y}_{i})\}_{i=1}^{N}$ is evaluated on a new set of input measurements $\bm{x}$, generated from the ground truth data generating process by linearly varying the position of the load $x_{\mathrm{F}}$. It can be seen that the effect of this variation on the measured response is largely captured by the data-driven component of the decoder, which overrides the known physics, despite the fact that the physics-based model includes the load position as an input parameter. The extent to which the data-driven components override the known physics can be inconsistent and hard to predict, and will depend on the neural network architectures and the physics of the problem at hand.
@@
The results presented in Figure 2 are obtained under the assumption of a factorized variational posterior. Ideally, a single encoder with shared parameters $q_{\mathrm{\phi}}(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}}|\bm{x})$ would be used for the three subsets of the latent space $\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. However, using a shared encoder leads to further degradation of the performance as noted by [34]. This example demonstrates how the interaction between the physics-based and data-driven components, which determines the resulting learned representation, depends on the capacity and flexibility of the individual machine learning components. Standard VAE offer no mechanism to control this interaction, and are therefore unable to guarantee that the physics will be utilized correctly in the presence of flexible NN-based decoder components.